OpenStudy (anonymous):

Taylor Series question Let f(x) = ln(1+x) Find a formula for the nth Taylor Polynomial of f, and then calculate p_n(1). Ok so I got up to the second derivative. Then for the third derivate, my response was 2/((1+x)^3) Why is the third derivative 2!/((1+x)^3).. where does the factorial come from?

OpenStudy (tkhunny):

Hard way! \(f(x) = \ln(1+x)\) \(f'(x) = \dfrac{1}{1+x} = 1 - x + x^{2} - x^{3} + ...\) \(f(x) = x - \dfrac{x^{2}}{2} + \dfrac{x^{3}}{3} - \dfrac{x^{4}}{4} + ...\) The funny factorial in the numerator is just to offset almost all of the factorial you would normally find in the denominator.

OpenStudy (anonymous):

i see.. so when we take the the third derivative.. instead of writing 6/((1+x)^4) we write 3!.. so when solving for p_n(x) it cancels out.. got it.. thank you!

OpenStudy (tkhunny):

Good work. You just have to build it.

OpenStudy (anonymous):

Thank you tkhunny!!